OpenStudy (anonymous):

the ratio of the number of hard cover books to the number of paper back books sold by a bookstore yesterday was 12 to 6. if the total number of hard cover books and paper back books sold by this bookstore was 204, how many paper back books were sold yesterday?"

OpenStudy (anonymous):

h/p = 12/6 h + p = 204

OpenStudy (anonymous):

the what do you do

OpenStudy (anonymous):

i meant then what do you do

OpenStudy (anonymous):

h/p = 2 h = 2p plug it in the second equation

OpenStudy (anonymous):

so h +2p=204 h+p=102

OpenStudy (anonymous):

then what doi do

OpenStudy (anonymous):

no is it 2p +p=204 3p=204 p=68

OpenStudy (anonymous):

sub here h + p = 204

OpenStudy (anonymous):

h = 2p

OpenStudy (anonymous):

hhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhhheeeeeeeeeeelllllllllppppppppppppp

OpenStudy (anonymous):

2p + p = 204 p = 204 /3 = 68

OpenStudy (anonymous):

hm..u got the ans and asking for help..))

OpenStudy (anonymous):

ya because the answers choices dont have 68 thats why

OpenStudy (anonymous):

h = 2p h = 68 * 2 = 136

OpenStudy (anonymous):

do anyone you guys have the kaplan shsat book that has over 1000 pages

OpenStudy (anonymous):

u need to find the No of pages

OpenStudy (mathteacher1729):

The key here is to find the link between the English sentences and the math. First, let's see if we can make sense of this... " the ratio of the number of hard cover books to the number of paper back books sold by a bookstore yesterday was 12 to 6. " This means that for every 12 hardcover books there were 6 paperbacks. You could neatly arrange each pile of hardcover and paperbacks like this: HHHHHH HHHHHH PPP PPP (notice how there are 12 Hs and 6 Ps.) But we can make it even simpler. for every ONE P there are THREE Hs. So we can split things up like this HHH P So for every 1 P there are three H. Does this make sense so far? Just in terms of interpreting that first sentence?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i need the answers the answer choice are 34 51 60 85 144

OpenStudy (mathteacher1729):

What part did you find confusing?

OpenStudy (anonymous):

hello does anyone out there have the Kaplan shsat book that has over 1000 pages

OpenStudy (anonymous):

the part that the answer isnt in the choices

OpenStudy (anonymous):

helllo anyon there

OpenStudy (mathteacher1729):

@Tahmina are you interested in learning how to solve these problems on your own or do you just want the answers to problems (A, B, C, D, etc)?

OpenStudy (anonymous):

no i want to learn but i dont have time

OpenStudy (anonymous):

you see my dilemma

OpenStudy (mathteacher1729):

Well, learning takes time. There's really no way around it. :-p I'm 29 and still learning.

OpenStudy (anonymous):

i need help badly im running out of time

OpenStudy (mathteacher1729):

I am glad to help walk you through the problem, but I do not give answers without having the student try and learn. That is my personal policy.

OpenStudy (anonymous):

well i tried

OpenStudy (anonymous):

i didnt have internet for the past week and then was a holiday so there went a day after that i tried this problem evryday

OpenStudy (anonymous):

anyone ther